If the black dot represents the wrist during a biceps curl, i.e., forearm rotation at the elbow. What is the linear velocity of the wrist as it travels from point 1 to 2, if ∆θ = 60°, time = 0.5 s, and length of the forearm (radius) = 0.5 m? (Conversion: 57.3⁰ = 1 rad)
Answer #1

I 0.5 Angular velocity w = 90 , 6o * 11/180 = 1.04719 t 0.5 S = 2.094395102 rad/s Linear velocity e-wr=2094395102 x 0.5 m/s =